OpenStudy (anonymous):

In the country of Glup, 1 glop is 3 glips and 4 glips are 5 globs. How many globs are 2 glops?

OpenStudy (anonymous):

who makes up these questions? -.-

OpenStudy (anonymous):

1 glop = 3 glips 4 glips = 5 globs # of globs = 2 glops

OpenStudy (anonymous):

8.75 glops..im pretty sure, @moneybird what did you get?

OpenStudy (anonymous):

i meant globs

OpenStudy (anonymous):

1/3 glops = 1 glip * 4 * 4 (Multiply four for the whole equation) then 4/3 glops = 4 glips which is also equal to 5 globs 4/3 glops = 5 globs *6/4 * 6 / 4 2 glops = 30 / 4 globs = 7.5 globs?